OpenStudy (anonymous):

Is the answer 1/4 ? The limit as x approaches pi/2 of (1-sinx) / (1+cos2x)

OpenStudy (luigi0210):

Is this your question: \[\lim_{x \rightarrow \frac{\pi}{2}}\frac{1-sinx}{1+\cos^2x}\]

OpenStudy (anonymous):

Could also be \[\cos 2x\]

OpenStudy (anonymous):

the bottom is 1+cos(2x)

OpenStudy (luigi0210):

Have you tried plugging in values?

OpenStudy (anonymous):

I tried but it says to use l'Hopital's Rule so I did and I got 1/4 I'm not sure if its right though

OpenStudy (luigi0210):

Wouldn't it be -1/4?

OpenStudy (luigi0210):

Because the derivative of cos2x is -2sin2x

OpenStudy (anonymous):

yeah and then i took the derivative again and got 4cos(pi) because when you plug in pi/2 you get 2pi/2 so the 2's cancel out don't they ?

OpenStudy (luigi0210):

When using L'hopitals don't we just take the derivative once?

OpenStudy (anonymous):

No we can take it more than once if we have to. Well, at least that's what my teacher told us

OpenStudy (luigi0210):

No, it's actually just once.

OpenStudy (anonymous):

It is ???

OpenStudy (luigi0210):

Wait, nevermind sorry, you were right

OpenStudy (luigi0210):

Sorry for wasting your time, but yea your answer is right :)

OpenStudy (anonymous):

So is the answer 1/4 or -1/4

OpenStudy (luigi0210):

1/4

OpenStudy (anonymous):

Okay thank you very much for your help :)

OpenStudy (luigi0210):

Yup :)